# Normalizing Flows are Capable Models for RL

### v2

@@ 1 Introduction
Normalizing Flows (NFs) are among the most flexible probabilistic models: they $(1)$ support both likelihood maximization and variational inference training, and $(2)$ enable efficient sampling and exact likelihood computation (see Table 1). Yet, NFs have received far less attention from the RL community, perhaps due to the (mis)conception that they have restricted architectures or training instabilities [64]. This paper demonstrates that neither of these notions hold NFs back in practice: NFs can match the expressivity of diffusion and autoregressive models while being significantly simpler. Moreover, unlike other models, NFs do not require additional machinery to be integrated with existing RL frameworks, enabling simpler algorithms that are competitive with if not better than their counterparts.
Our primary contribution is to shed light on the effectiveness and NFs in RL. Building on prior work [45, 19], we develop a simple NF architecture and evaluate it across 82 tasks spanning five different RL settings, including offline imitation learning, offline RL, goal-conditioned RL, and unsupervised RL:
@@ 2 Related Work
This work builds on the rich literature of normalizing flow architectures [18, 72, 45, 46, 39, 65, 89]. NICE [18] introduce the idea of stacking neural function blocks that perform shifting, making density estimation efficient. RealNVP [19] extends NICE to non volume preserving layers that also perform a scaling operation. GLOW [45] further propose applying an invertible generalized permutation and a special normalization layer (ActNorm) to improve expressivity.
@@ 5 A simple and efficient NF architecture for general RL
Normalization.
While previous approaches used special variants of normalization layers ([19], [45]), we find that just using LayerNorm [3] in the coupling network’s weights was sufficient of training deep NFs.
@@ 6 Experiments
In all experiments, we use the architecture described in Section 5 to implement all algorithms described in Section 4. Implementation details for all algorithms are provided in Appendix A. Across all problems (IL, offline RL, online GCRL), we present results on a total of 82 tasks (see Fig. 2 for visualization of tasks) with 5 seeds each, which show that NFs are generally capable models that can be effectively applied to a diverse set of RL problems. Unless stated otherwise, all error bars represent $\pm$ standard deviation across seeds.
@@ 7 Conclusion
The authors are pleased to acknowledge that the work reported on in this paper was substantially performed using the Princeton Research Computing resources at Princeton University which is consortium of groups led by the Princeton Institute for Computational Science and Engineering (PICSciE) and Office of Information Technology’s Research Computing. The authors thank Emmanuel Bengio, Liv d’Aliberti and Eva Yi Xie for their detailed reviews on our initial drafts. The authors thank Seohong Park for providing code for some baselines and Catherine Ji for helping find a bug in our code. The authors also thank the members of Princeton RL lab for helpful discussions.

### v3

@@ 1 Introduction
Normalizing Flows (NFs) are among the most flexible probabilistic models: they $(1)$ support both likelihood maximization and variational inference training, and $(2)$ enable efficient sampling and exact likelihood computation (see Table 1). Yet, NFs have received far less attention from the RL community, perhaps due to the (mis)conception that they have restricted architectures or training instabilities [64]. This paper demonstrates that neither of these notions hold NFs back in practice: NFs can match the expressivity of diffusion and autoregressive models while being significantly simpler. Moreover, unlike other models, NFs do not require additional machinery to be integrated with existing RL frameworks, enabling simpler algorithms that are competitive with, if not better than their counterparts.
Our primary contribution is to shed light on the effectiveness and NFs in RL. Building on prior work [45, 19, 89], we develop a simple NF architecture and evaluate it across 82 tasks spanning five different RL settings, including offline imitation learning, offline RL, goal-conditioned RL, and unsupervised RL:
@@ 2 Related Work
This work builds on the rich literature of normalizing flow architectures [18, 72, 45, 46, 39, 65, 89]. NICE [18] introduce the idea of stacking neural function blocks that perform shifting, making density estimation efficient. RealNVP [19] extends NICE to non volume preserving layers that also perform a scaling operation. GLOW [45] further propose applying an invertible generalized permutation and a special normalization layer (ActNorm) to improve expressivity. Recently, Zhai et al., [89] propose a scalable NF architecture and a sampling technique to achieve SOTA results on vision benchmarks.
@@ 5 A simple and efficient NF architecture for general RL
Other details. While previous approaches used special variants of normalization layers ([19, 45]), we find that just using LayerNorm [3] in the coupling network’s weights was sufficient of training deep NFs. For NFs trained purely using MLE (Section 6.1, Section 6.2), we found the noising and denoising technique proposed by [89] (see Section A.1 for details) to be crucial for good performance.
@@ 6 Experiments
In all experiments, we use the architecture described in Section 5 to implement all algorithms from Section 4, with further implementation details in Appendix A. Across all problems (IL, offline RL, online GCRL), we present results on a total of 82 tasks (see Fig. 2) with 5 seeds each, showing that NFs are generally capable models applicable to diverse RL problems. Unless stated otherwise, all error bars represent $\pm$ standard deviation across seeds.
@@ 7 Conclusion
The authors are pleased to acknowledge that the work reported on in this paper was substantially performed using the Princeton Research Computing resources at Princeton University which is consortium of groups led by the Princeton Institute for Computational Science and Engineering (PICSciE) and Office of Information Technology’s Research Computing. The authors thank Emmanuel Bengio, Liv d’Aliberti and Eva Yi Xie for their detailed reviews on our initial drafts. The authors thank Seohong Park for providing code for some baselines and Catherine Ji for helping find a bug in our code. The authors also thank Shuangfei Zhai and the members of Princeton RL lab for helpful discussions.